# NeRF-Supervised Feature Point Detection and Description

### v2

@@ 1 Introduction
This paper aims at leveraging image synthesis with neural radiance fields (NeRFs) as a more realistic way of generating multi-view training data for feature detection and description models (see Fig. 1). Since NeRFs require multi-view data to synthesise novel views, we can no longer rely on the single image datasets typically used for training the above-mentioned homography-based methods. Therefore, we create our own dataset consisting of indoor and outdoor image sequences around static scenes and reconstruct all of them with NeRFacto [13]. This enables the generation of arbitrary viewpoints of each scene consistent with a pinhole projective model with known point-to-point mappings via point re-projection. We propose a general methodology to upgrade state-of-the-art homography-based methods to train them on projective views synthesised from NeRF-type algorithms. Our contributions are as follows:
  • We create a new multi-view dataset consisting of images from 10 different indoor and outdoor scenes, and a total of 10000 NeRF-synthesised views from these scenes with corresponding depth maps, and intrinsic and extrinsic parameters.
@@
  • We re-train adapted versions of SuperPoint [6] and SiLK [14] using our NeRF-synthesised data and compare them against the original baselines trained on the much larger MS-COCO dataset. We outperform the original baselines for relative pose estimation on ScanNet, YFCC100M and MegaDepth datasets [15, 16, 17], with similar performance for pair-wise point cloud registration, while only slightly under-performing on the HPatches homography estimation benchmark.
@@
Figure 1: Visual representation of multi-view data. Learning-based detectors and descriptors achieve supervision on single-view datasets by simulating different viewpoints through homographic warpings to the input image $I$ (see Fig. 1(a)) resulting in $I^{{}^{\prime}}$ (see Fig. 1(b)). However, we achieve supervision by directly sampling a NeRF-rendered image from a different viewpoint (see Fig. 1(c)).
@@ 3.1 NeRF Dataset
We build a NeRF dataset for 10 indoor and outdoor scenes, containing 10000 synthetic images with their corresponding intrinsic/extrinsic parameters and depth maps. We note that this dataset is extremely small compared to what is typically used to train state-of-the-art feature detection (30 times smaller than MS-COCO, 250 times smaller than ScanNet). Our goal is to show that leveraging NeRF supervision enables state-of-the-art performance on multiple benchmarks with significantly less training data.
We captured four indoor and one outdoor scenes with an iPhone 10 at 4K resolution, and one indoor and one outdoor scene with a Samsung A52 at 4k resolution. We also utilise two blender-generated indoor scenes with images sourced from [33] and one outdoor scene publicly available through [34] as they are open-source scenes that consist of high-quality RGB images. While different training sets could be constructed using existing publicly available data, we focus on a trade-off between scene diversity and minimal dataset size.
We obtain camera poses for all images using COLMAP [35] and use them to generate synthetic views. We trained a Nerfacto model for each scene, which is part of the NeRFStudio framework [13] and merges improvements from a range of NeRF implementations [36, 34, 37, 38]. For each scene, we rendered 1000 synthetic views and corresponding depth maps with $44^{\circ}$ field of view and $640\times 480$ resolution. We use NeRFStudio’s interactive real-time viewer to define custom camera trajectories for each scene that ensure diversity in viewpoints while maintaining global scene overlap. Trajectories consist of uniformly sampled smooth trajectories around the scene coupled with small rotations around the camera centre.
@@ 3.2 NeRF Point Re-Projection
Assuming that reconstructed depths $d$ are accurate, the above equations enable supervising a point detector, by designing a loss function that promotes a set of re-projected points $\mathbf{p}^{\prime}$ detected in view $I$ to align with points independently detected on view $I^{\prime}$. However, many distinctive interest points lie on scene edges at a depth discontinuity. This means that even a small deviation in $\mathbf{p}$ can cause significant variations in $d$, corresponding to the depths of the background and foreground around the edges. To ensure stable re-projection around edges, we differentiate between foreground and background by focusing only on the foreground around the edges, excluding the background. To achieve this, we compute $d$ within a $5\times 5$ pixel window centred around $\mathbf{p}$. If the difference between the maximum and minimum depths within this patch does not exceed a predefined threshold $\epsilon_{d}$, we take $d$ at the interest point position. Otherwise, we use the minimum depth within the window (see example in Fig. 2).
@@
Figure 2: Depth window estimation. The interest points, depicted in red and purple situated at the painting’s frame in Fig. 2(a), exhibit misprojection onto image $I^{{}^{\prime}}$ without the depth window estimation method. However, misprojection errors are effectively mitigated by utilising the depth window estimation method as seen in Fig. 2(c).
@@ 4.1 SiLK-PrP
In the original SiLK approach, image pairs with ground-truth bi-directional correspondences are generated on the fly during training with homography warpings. We rely on our pre-rendered NeRF dataset described in section 3.1, to avoid significant computational overhead caused by on-the-fly image rendering. During training, we randomly sample image pairs $I,I^{{}^{\prime}}$ from the same scene, Eqs. 1 and 2 are then utilised during training to generate dense ground-truth point correspondences, discretised at the pixel level.
@@ 4.2 SuperPoint-PrP
SuperPoint is a self-supervised method that goes through multiple stages of training. Firstly, an encoder with a detector head is trained in a fully supervised manner to detect corners on a synthetic dataset with simple geometric shapes. The pre-trained detector, called MagicPoint, is used to generate pseudo-ground truth interest points on a real dataset (e.g. MS-COCO) using a process called homographic adaptation. This involves generating several homography warped copies of a training image, passing them through the trained detector, unwarping all predictions back to the original training image, and finally aggregating all unwarped predictions to generate the pseudo label. This process can be repeated several times, where after each round of homographic adaptation, the detector head is further trained, updating the pseudo-ground truth labels, while improving the detector head’s generalisability. After a final round of homographic adaptation, a descriptor head is added to the model, which is jointly trained with the whole model with an additional descriptor loss term on the real dataset resulting in the SuperPoint model.
With the NeRF training data approach, we use the same architecture as the original SuperPoint model. However, modify the homographic adaptation process as well as the descriptor loss to handle our NeRF training data and Point Re-Projection process.
@@ 4.2.1 Projective Adaptation
Similarly to our SiLK-PrP approach, we consider pre-rendered NeRF images as training data. In this case, we exploit the fact that training sequences for each scene are generated along a continuous trajectory.
We randomly sample from a window of 20 consecutive rendered NeRF images $I_{i\rightarrow i+19}$. We take pseudo-ground-truth points for the initial image $I_{i}$ and randomly sample a set of 14 of the remaining 19 images as the warped versions of $I_{i}$. An image from the randomly sampled set is denoted as $I_{r}$.
The probability heatmap $\mathbf{{H}_{i}}$ for image $I_{i}$ is then computed. For each image $I_{r}$ in the randomly selected set, its corresponding probability heatmap $\mathbf{{H}_{r}}$ is computed. Interest points $\mathbf{{p}_{r}}$ are then extracted from each heatmap, employing non-maximum suppression to filter out feature points that are too close to each other.
Once interest points $\mathbf{{p}_{r}}$ are obtained, the PrP process is applied to project the interest points aligning them with the initial frame ${I}_{i}$, to clarify the projected interest points from $I_{r}$ to $I_{i}$ can be denoted as $\mathbf{{p}^{{}^{\prime}}_{i}}$.
Having obtained the projected interest points $\mathbf{{p}^{{}^{\prime}}_{i}}$, we generate a mask denoted as $\mathbf{{H}_{i^{\prime}}}$, for each interest point in $\mathbf{{p}^{{}^{\prime}}_{i}}$ coordinates on the binary mask, a 3x3 patch centred around the interest point before projection is extracted from the probability heatmap $\mathbf{{H}_{r}}$, the patch is then applied on the binary heatmap $\mathbf{{H}_{i^{\prime}}}$. Thus, mask $\mathbf{{H}_{i^{\prime}}}$, serves as the probability heatmap of rendered image $I_{r}$ projected onto the input render frame ${I}_{i}$.
Finally, the input’s rendered image probability heatmap $\mathbf{{H}_{i}}$ and mask $\mathbf{{H}_{i^{\prime}}}$ for each frame, are aggregated together, non-maximum suppression is applied once more on the final aggregated heatmap to obtain pseudo-ground-truth interest points for the input rendered image $I_{i}$.
@@ 4.2.2 Descriptor Loss
where $m_{p}$ is the hinge loss’s positive margin, $m_{n}$ is the negative margin and $\lambda_{d}$ balances between the positive and negative correspondences. Additionally, $S$ in Eq. 3 is a homography-induced correspondence, defined as:
@@
We reformulate the homography-induced correspondence of SuperPoint [6], in such a way that centre pixel $p_{hw}$ can be transformed using the PrP process rather than transforming it through homography $\mathcal{H}$. Let $\mathcal{C}(\cdot)$ represent a function that takes $p_{hw}$ as an input, along with all parameters utilised in the PrP process (see Eqs. 1 and 2), for simplicity, these parameters can be combined and denoted as $M$. Therefore, the PrP-induced correspondence can be defined as:
@@ 4.3 Implementation details
For all experiments, we present results using our training pipeline for both SuperPoint-PrP and SiLK-PrP. Additionally, both SuperPoint-PrP and SiLK-PrP were trained exclusively on the NeRF dataset. Furthermore, we report results for the baseline SuperPoint[6] model using our own reproduced implementation, which closely follows the training hyperparameters of the public version of SuperPoint [39].
To mitigate misprojection errors during the PrP process, we employ a random sampling approach for selecting image $I^{{}^{\prime}}$ from a subset of the training data. This subset is constrained within a range determined by a lower limit, $\lambda_{l}$, and an upper limit, $\lambda_{u}$, based on the input image $I$. Specifically, we set $\lambda_{l}$ to 70 and $\lambda_{u}$ to 150, ensuring a balance between viewpoint changes while avoiding projection errors. Moreover, the depth window estimation process threshold $\epsilon_{d}$ was set as $3e^{-2}$ meters.
Additionally, for SuperPoint-PrP and SiLK-PrP training, we apply the same photometric augmentation applied on the baseline models [6, 14] such as random brightness and contrast, Gaussian noise, motion blur, but do not apply homographic augmentation such as rotation, scaling or translation. Lastly, all models are trained using PyTorch [40] and a single NVIDIA RTX 3090 Ti GPU.
@@
Similar to SuperPoint’s baseline training pipeline, we obtain the base MagicPoint model by conducting training on the Synthetic Shapes dataset for 200,000 iterations. The MagicPoint model is identical to the SuperPoint model omitting the descriptor head, while the Synthetic Shapes dataset consists of synthetic 2D geometrical shapes such as lines, triangles and cubes.
To produce the MagicPoint-PrP model, we further train the base MagicPoint model through two rounds of Projective Adaptation on the NeRF dataset, each lasting 30,000 iterations. Finally, the SuperPoint-PrP model is obtained by training both the detector and descriptor heads of MagicPoint-PrP once again on the NeRF dataset for a total of 300,000 iterations. The training process is performed using the baseline SuperPoint model hyperparameters with ADAM optimiser, a learning rate of $1e^{-3}$ and betas set to (0.9, 0.999). The batch size is set at 32 for the MagicPoint-PrP model and 2 for the SuperPoint-PrP model. Lastly, for the Baseline SuperPoint model, the default and optimum homography-induced correspondence threshold $\epsilon_{s}$ in Eq. 7 is 8. However, experiments showed that the optimum PrP-induced correspondence threshold $\epsilon_{s}$ in Eq. 10 is 4.
@@ 5 Experiments
In this section, we assess the performance of PrP-supervised models compared to their corresponding baselines. Evaluations include interest point detection quality and homography estimation on HPatches [41], indoor relative pose estimation and point cloud registration on ScanNet [15], and outdoor relative pose estimation on YFCC100M and MegaDepth [16, 17]. For all evaluations, we use mutual nearest neighbour for feature matching. Through these experiments, we analyse the following:
  • Our NeRF dataset is approximately $\frac{1}{30}$ the size of the MS-COCO dataset [20], and $\frac{1}{250}$ the size of ScanNet’s [15]. Despite the relatively small scale of our training dataset, we analyse the generalisability of the NeRF-trained models on real-world data and various conditions across multiple datasets. Surprisingly, the PrP models trained on our NeRF datasets consistently demonstrate comparable results in most evaluations. This raises the question of the necessity for large open-source datasets in training learning-based detectors and descriptors.
@@
  • As we explicitly render each scene in our NeRF dataset from various viewpoints, simulating real camera motion instead of relying on homography to mimic camera motions, we explore whether models supervised by the PrP process demonstrate improvement in their relative pose estimation capabilities. Additionally, we assess whether PrP-supervised models require fewer iterations during training compared to the homography-supervised baselines. Results show that not only do the PrP-supervised models exhibit enhancement in relative pose estimation but also converge faster (i.e. SiLK-PrP converged in approximately 27K iterations compared to 50K for SiLK).
@@ 5.1 Homography Estimation
We follow [24, 14] and resize the shorter image side to a dimension of 480. To evaluate interest point detection quality and performance, we report the repeatability metric as in [6, 14]. Moreover, we evaluate the keypoint descriptors, we report the mean matching accuracy (MMA) i.e. the ratio of matches where reprojection error is below a specified threshold [21], and matching score i.e. The ratio between correctly matched points and a total number of detected points. Lastly, we use OpenCV’s RANSAC algorithm to estimate the homography between an image pair and report the homography estimation accuracy and area under the curve (AUC) error over all image pairs. We keep the top 1k detected interest points for SuperPoint and 10k for SiLK.
@@
We assess SuperPoint-PrP by comparing it to the baseline SuperPoint model [6]. Additionally, we introduce a hybrid variant of SuperPoint-PrP denoted as "SP-PrP-Hyb". This variant incorporates the MagicPoint model, trained through two rounds of Homographic Adaptation, followed by training the SuperPoint model on the NeRF dataset using our PrP process and Projective Adaptation. Furthermore, we present two versions of PrP SiLK models; the first, SiLK-PrP, is trained without rotation or scaling augmentation as detailed in Sec. 4.1. The second, SiLK-PrP-Aug, includes rotation and scaling augmentations during training.
@@
As seen in Tab. 1, the PrP-supervised models are outperformed by their baseline counterparts [14, 6] in all of the HPatches metrics. Nonetheless, this outcome is expected as both SiLK-PrP and SuperPoint-PrP are not rotation or scale-invariant, in contrast to their corresponding baseline models. On the other hand, despite SiLK-PrP-Aug’s efforts to incorporate in-place augmentation for rotation and scale invariance (see Fig. 3), it falls short compared to the SiLK model.
Tab. 2 further delves into SiLK-PrP-Aug’s homography estimation performance, revealing its homography estimation aligns with SiLK in scenes with varying viewpoints. However, a dip in performance is observed in scenes with varying illumination conditions. This discrepancy contributes to the observed performance gap between SiLK-PrP-Aug and SiLK [14] in Tab. 1. We opted to use the same photometric augmentation used for SiLK [14] when training SiLK-PrP and SiLK-PrP-Aug, however, optimising photometric augmentation parameters during training SiLK-PrP and SiLK-PrP-Aug on the NeRF dataset might offer improved results.
@@
Table 2: HPatches metrics separated. SiLK-PrP and SiLK are similar in estimating homography for scenes with varying viewpoints, while SiLK-PrP falls behind in varying illumination scenes.
@@
Figure 3: Rotation and Scale invariance. In Fig. 3(b), it is evident that SiLK-PrP lacks rotation and scale invariance compared to SiLK (see Fig. 3(a)) [14]. Incorporating rotation and scaling SiLK-PrP’s training process, rotation and scale invariance is achieved (see Fig. 3(c)).

### v3

@@ 1 Introduction
This paper aims at leveraging image synthesis with neural radiance fields (NeRFs) as a more realistic way of generating multi-view training data for feature detection and description models (cf. Fig. 1). Since NeRFs require multi-view data to synthesise novel views, we can no longer rely on the single image datasets typically used for training the above-mentioned homography-based methods. Therefore, we create our dataset consisting of indoor and outdoor image sequences around static scenes and reconstruct all of them with NeRFacto [13]. This enables the generation of arbitrary viewpoints of each scene consistent with a pinhole projective model with known point-to-point mappings via point re-projection. We propose a general methodology to upgrade state-of-the-art homography-based methods to train them on projective views synthesised from NeRF-type algorithms. Our contributions are as follows:
  • We create a new multi-view dataset consisting of images from 10 different indoor and outdoor scenes, and a total of 10,000 NeRF-synthesised views from these scenes with corresponding depth maps, and intrinsic and extrinsic parameters.
@@
  • We re-train adapted versions of SuperPoint [6] and SiLK [14] using our NeRF-synthesised data and compare them against the original baselines trained on the much larger MS-COCO dataset. We outperform the original baselines for relative pose estimation on ScanNet, YFCC100M and MegaDepth datasets [15, 16, 17], with similar performance for pair-wise point cloud registration, while only slightly underperforming on the HPatches homography estimation benchmark. Our code is available at SiLK-PrP and SuperPoint-PrP.
@@
Figure 1: Visual representation of multi-view data. Learning-based detectors and descriptors achieve supervision on single-view datasets by simulating different viewpoints through homographic warpings to the input image $I$ (cf. Fig. 1(a)) resulting in $I^{{}^{\prime}}$ (cf. Fig. 1(b)). However, we achieve supervision by directly sampling a NeRF-rendered image from a different viewpoint (cf. Fig. 1(c)).
@@ 3.1 NeRF Dataset
We build a NeRF dataset for 10 indoor and outdoor scenes, containing 10,000 synthetic images with their corresponding intrinsic/extrinsic parameters and depth maps. We note that this dataset is extremely small compared to what is typically used to train state-of-the-art feature detection (30 times smaller than MS-COCO, 250 times smaller than ScanNet). Our goal is to show that leveraging NeRF supervision enables state-of-the-art performance on multiple benchmarks with significantly less training data.
We captured four indoor and one outdoor scene with an iPhone 10 at 4K resolution, and one indoor and one outdoor scene with a Samsung Galaxy A52 at 4k resolution. We also utilise two blender-generated indoor scenes with images sourced from [33] and one outdoor scene publicly available through [34] as they are open-source scenes that consist of high-quality RGB images. While different training sets could be constructed using existing publicly available data, we focus on a trade-off between scene diversity and minimal dataset size.
We obtain camera poses for all images using COLMAP [35] and use them along with the images to train a NeRF model for synthesising synthetic views. We trained a Nerfacto model for each scene, which is part of the NeRFStudio framework [13] and merges improvements from a range of NeRF implementations [36, 34, 37, 38]. For each scene, we rendered 1000 synthetic views and corresponding depth maps with $44^{\circ}$ field of view and $640\times 480$ resolution. We use NeRFStudio’s interactive real-time viewer to define custom camera trajectories for each scene that ensure diversity in viewpoints while maintaining global scene overlap. Trajectories consist of uniformly sampled smooth trajectories around the scene coupled with small rotations around the camera centre.
@@ 3.2 NeRF Point Re-Projection
Assuming that reconstructed depths $d$ are accurate, the above equations enable supervising a point detector, by designing a loss function that promotes a set of re-projected points $\mathbf{p}^{\prime}$ detected in view $I$ to align with points independently detected on view $I^{\prime}$. However, many distinctive interest points lie on scene edges at a depth discontinuity. This means that even a small deviation in $\mathbf{p}$ can cause significant variations in $d$, corresponding to the depths of the background and foreground around the edges. To ensure stable re-projection around edges, we differentiate between foreground and background by focusing only on the foreground around the edges, excluding the background. To achieve this, we compute $d$ within a $5\times 5$ pixel window centred around $\mathbf{p}$. If the difference between the maximum and minimum depths within this patch does not exceed a predefined threshold $\epsilon_{d}$, we take $d$ at the interest point position. Otherwise, we use the minimum depth within the window (cf. Fig. 2).
@@
Figure 2: Depth window estimation. The interest points, depicted in red and purple situated at the painting’s frame edge in Fig. 2(a), exhibit misprojection onto image $I^{{}^{\prime}}$ without the depth window estimation method. However, misprojection errors are effectively mitigated by utilising the depth window estimation method as seen in Fig. 2(c).
@@ 4.1 SiLK-PrP
In the original SiLK approach, image pairs with ground-truth bi-directional correspondences are generated on the fly during training with homography warpings. We rely on our pre-rendered NeRF dataset described in Sec. 3.1, to avoid significant computational overhead caused by on-the-fly image rendering. During training, we randomly sample image pairs $I,I^{{}^{\prime}}$ from the same scene, Eqs. (1) and (2) are then utilised during training to generate dense ground-truth point correspondences, discretised at the pixel level.
@@ 4.2 SuperPoint-PrP
SuperPoint is a self-supervised method that goes through multiple stages of training. Firstly, an encoder with a detector head is trained in a fully supervised manner to detect corners on a synthetic dataset with simple geometric shapes. The pre-trained detector, called MagicPoint, is used to generate pseudo-ground truth interest points on a real dataset (e.g. MS-COCO) using a process called homographic adaptation. This involves generating several homography warped copies of a training image, passing them through the trained detector, unwarping all predictions back to the original training image, and finally aggregating all unwarped predictions to generate the pseudo labels. This process can be repeated several times, where after each round of homographic adaptation, the detector head is further trained, updating the pseudo-ground truth labels, while improving the detector head’s generalisability. After a final round of homographic adaptation, a descriptor head is added to the model, which is jointly trained with the whole model with an additional descriptor loss term on the real dataset resulting in the SuperPoint model.
With the NeRF training data, we retain the original SuperPoint architecture but adapt the homographic adaptation process and descriptor loss to handle our NeRF training data and the Point Re-Projection process.
@@ 4.2.1 Projective Adaptation
We propose a novel method for generating pseudo-ground-truth labels using the PrP process instead of homography formulations. Our approach takes advantage of the continuous trajectory used to render the NeRF images.
The process begins by randomly sampling a window of 20 consecutive rendered NeRF images, denoted as $I_{i\rightarrow i+19}$. We designate the initial image $I_{i}$ as our reference and randomly select 14 of the remaining 19 images as warped versions of $I_{i}$. We refer to any image from this randomly sampled set as $I_{r}$. For both $I_{i}$ and each $I_{r}$, we compute probability heatmaps $\mathbf{H}_{i}$ and $\mathbf{H}_{r}$ respectively. We then extract interest points $\mathbf{{p}_{r}}$ from each $\mathbf{H}_{r}$ using non-maximum suppression to ensure sufficient spacing between feature points.
Next, we apply the PrP process to project the interest points $\mathbf{p}_{r}$ from each image $I_{r}$ onto $I_{i}$, resulting in projected points $\mathbf{p}_{i^{\prime}}$. For all points in $\mathbf{p}_{i^{\prime}}$, we create a mask $\mathbf{H}_{i^{\prime}}$ by extracting a 3x3 patch centred on the original (unprojected) point location from $\mathbf{H}_{r}$ and applying it to $\mathbf{H}_{i^{\prime}}$ at the projected point location. Therefore, $\mathbf{H}_{i^{\prime}}$ effectively represents the probability heatmap of $I_{r}$ projected onto $I_{i}$.
Finally, we aggregate $\mathbf{H}_{i}$ with all $\mathbf{H}_{i^{\prime}}$ masks and apply non-maximum suppression to the aggregated heatmaps to obtain the final set of pseudo-ground-truth interest points for $I_{i}$. This approach leverages the continuous nature of NeRF-rendered sequences to create pseudo-ground-truth labels without relying on homography approximations.
@@ 4.2.2 Descriptor Loss
where $m_{p}$ is the hinge loss’s positive margin, $m_{n}$ is the negative margin and $\lambda_{d}$ balances between the positive and negative correspondences. Additionally, $S$ in Eq. (3) is a homography-induced correspondence, defined as:
@@
We reformulate the homography-induced correspondence of SuperPoint [6], in such a way that centre pixel $p_{hw}$ can be transformed using the PrP process rather than transforming it through homography $\mathcal{H}$. Let $\mathcal{C}(\cdot)$ represent a function that takes $p_{hw}$ as an input, along with all parameters utilised in the PrP process (cf. Eq. (1) and Eq. (2)), for simplicity, these parameters can be combined and denoted as $M$. Therefore, the PrP-induced correspondence can be defined as:
@@ 4.3 Implementation details
For all experiments, we present results for SuperPoint-PrP and SiLK-PrP using our training pipeline, both trained exclusively on the NeRF dataset. Additionally, we report the baseline SuperPoint results based on our implementation, which closely follows the public version’s training hyperparameters [39].
To mitigate misprojection errors during the PrP process, we randomly sample image $I^{{}^{\prime}}$ from a subset of the training data. This subset is constrained by a lower limit of $\lambda_{l}=70$ and an upper limit of $\lambda_{u}=150$ relative to the input image $I$, ensuring a balance between viewpoint changes while avoiding projection errors. The depth window threshold $\epsilon_{d}$ described in Sec. 3.2 is set to $3e^{-2}$ meters.
During SuperPoint-PrP and SiLK-PrP training, we apply the same photometric augmentations as baseline models [6, 14], but omit homographic augmentations, such as rotation, scaling or translation. All models are trained using PyTorch [40] on a single NVIDIA RTX 3090 Ti GPU.
@@
Similar to the baseline training pipeline, we train MagicPoint (SuperPoint without descriptor head) on Synthetic Shapes (2D geometric shapes like lines, triangles, and cubes) for 200,000 iterations. We obtain MagicPoint-PrP, by training the MagicPoint model on the NeRF dataset for two rounds of Projective Adaptation, each for 30,000 iterations. The SuperPoint-PrP model is then obtained by training both detector and descriptor heads of MagicPoint-PrP on the NeRF for 300,000 iterations.
We use baseline SuperPoint hyperparameters with ADAM optimiser, a learning rate of $1e^{-3}$, and betas set to (0.9, 0.999). Batch sizes are 32 for MagicPoint-PrP and 2 for SuperPoint-PrP. For the baseline SuperPoint model, the default homography-induced correspondence threshold $\epsilon_{s}$ in Eq. (7) is 8. However, for SuperPoint-PrP, experiments showed that the optimum PrP-induced correspondence threshold $\epsilon_{s}$ in Eq. (10) is 4.
@@ 5 Experiments
This section assesses the PrP-supervised models against their baselines, evaluating detection quality and homography estimation on HPatches [41], indoor relative pose estimation and point cloud registration on ScanNet [15], and outdoor relative pose estimation on YFCC100M and MegaDepth [16, 17]. All evaluations use mutual nearest neighbour matching. Through these experiments, we analyse:
  • The NeRF dataset is about $\frac{1}{30}$ the size of MS-COCO [20] and $\frac{1}{250}$ of ScanNet [15]. Therefore, we evaluate the generalisability of NeRF-trained models on real-world data across various datasets. Surprisingly, the PrP models trained on the NeRF dataset perform comparably in most evaluations, questioning the need for large datasets to train learning-based detectors and descriptors.
@@
  • By rendering each NeRF scene from various viewpoints while simulating real camera motion, we evaluate whether PrP-supervised models demonstrate improvement in relative pose estimation, and require fewer training iterations than homography-based baselines. Results show that the PrP-supervised models not only exhibit enhancement in relative pose estimation but also converge faster (e.g. SiLK-PrP in 27K iterations vs. 50K for SiLK).
@@ 5.1 Homography Estimation
We follow prior work [24, 14] and resize the shorter image side to 480 pixels. We report the repeatability metric as in [6, 14] to evaluate interest point detection quality. We assess keypoint descriptors with the mean matching accuracy (MMA) - the ratio of matches with reprojection error below a threshold [21], and the matching score - the ratio of correctly matched points to total detected points. We use OpenCV’s RANSAC to estimate the homography between image pairs and report the homography estimation accuracy and area under the curve (AUC) error over all image pairs. We retain the top 1k detected points for SuperPoint and 10k for SiLK.
@@
We assess SuperPoint-PrP by comparing it to the baseline SuperPoint model [6]. We also introduce a hybrid variant, "SP-PrP-Hyb," where the MagicPoint model undergoes training on two rounds of Homographic Adaptation before training the SuperPoint model on the NeRF dataset using the PrP process and Projective Adaptation. To assess the impact of the dataset size, we include two additional variants: "SP-(10k COCO)" a version of the baseline model trained on 10k MS-COCO images matching the size of the NeRF dataset, and "SP-PrP-(2k)" a version of SuperPoint-PrP trained on 2k NeRF images (200 randomly sampled images from each scene) instead of the full NeRF dataset. Finally, we present two PrP SiLK models: "SiLK-PrP," trained without rotation or scaling augmentation as detailed in Sec. 4.3, and "SiLK-PrP-Aug," which includes the aforementioned augmentations during training.
@@
As seen in Tab. 1, the PrP-supervised models underperform compared to their baselines [14, 6] across all HPatches metrics. This is expected as both SuperPoint-PrP and SiLK-PrP are not rotation or scale-invariant. Despite SiLK-PrP-Aug’s efforts to incorporate in-place augmentation for rotation and scale invariance (cf. Fig. 3), it still lags behind SiLK [14]. The SP-(10k COCO) model shows a significant performance drop compared to the baseline SuperPoint [6], while SP-PrP-(2k) remains competitive with SP-PrP, highlighting the advantages of the multi-view NeRF dataset and the robustness of the PrP supervision.
Nonetheless, Tab. 2 reveals that SiLK-PrP-Aug’s homography estimation is consistent with SiLK [14] in varying viewpoints but drops in scenes with varying illumination. This discrepancy contributes to the observed performance gap between SiLK-PrP-Aug and SiLK [14] in Tab. 1. Although the same photometric augmentation as SiLK [14] was used for training SiLK-PrP and SiLK-PrP-Aug, optimising these parameters during training SiLK-PrP and SiLK-PrP-Aug on the NeRF dataset may yield improved homography estimation accuracy and greater robustness to changes in illumination.
@@
Table 2: HPatches metrics separated. SiLK-PrP and SiLK are similar in estimating homography for scenes with varying viewpoints, while SiLK-PrP falls behind in scenes with varying illumination.
@@
Figure 3: Rotation and Scale invariance. In Fig. 3(b), it is evident that SiLK-PrP lacks rotation and scale invariance compared to SiLK (cf. Fig. 3(a)) [14]. Incorporating rotation and scaling augmentations during SiLK-PrP’s training, rotation and scale invariance is achieved (cf. Fig 3(c)).